OpenStudy (anonymous):

5t-5/-10 =70 need a solution

jimthompson5910 (jim_thompson5910):

5t-5/-10 =70 5t(-10)-5 =70(-10) ... Multiply EVERY term by -10 to clear out the fraction -50t-5 = -700 -50t = -700+5 -50t = -695 t = -695/(-50) t = 139/10